Use the confidence level and sample data to find a confidence interval for estimating the population μ. Round your answer to the same number of decimal places as the sample mean.
39 packages are randomly selected from packages received by a parcel service. The sample has a mean weight of 29.9 pounds and a standard deviation of 3.0 pounds. What is the 95% confidence interval for the true mean weight, μ, of all packages received by the parcel service?

Answer by rothauserc(4718)

sample mean is 29.9, sample standard deviation is 3.0 pounds and sample size is 39
:
since sample size is > 30, we can use a normal distribution
:
alpha(a) = 1 - (95/100) = 0.05
:
critical probability(p*) = 1 - (a/2) = 0.975
:
critical value(CV) is the z-score associated with p*, the CV is 1.96
:
standard error(SE) = 3/square root(39) = 0.4804
:
margin of error(ME) = CV * SE = 1.96 * 0.4804 = 0.9416 approximately 0.9
:
*****************************************************************
95% confidence interval = 29.9 + or - 0.9, which is (29.0, 30.8)
*****************************************************************
: